OpenStudy (accidentalaichan):

Find A union B. A: {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12} B: {−2, −1, 1, 4, 7, 11}

OpenStudy (accidentalaichan):

Oh, I get it. Thanks! So, {-2, -2, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12}?

OpenStudy (accidentalaichan):

*-2, -1

OpenStudy (psymon):

Am I saying something incorrect?

OpenStudy (ankit042):

NO @Psymon fumbled with Union and intersection ....your previous answer looks correct..just don't repeat the terms

OpenStudy (psymon):

Ah, you're right, sorry x_x Mixed them up

OpenStudy (psymon):

Reversed the order. Geez, I mess up a lot o.o

OpenStudy (accidentalaichan):

It's quite alright. Thanks for helping @ankit042 and @Psymon

OpenStudy (anonymous):

\[\bf A \cup B = A+B-(A ∩ B)\]

OpenStudy (psymon):

Eh, thanks @AccidentalAiChan x_x I embarrass myself too much on here, though, lol x_x

OpenStudy (ankit042):

Haha It is Ok to make mistakes as far as you learn from them!

OpenStudy (anonymous):

@AccidentalAiChan Let's assume that a set \(\bf A = {a_1,a_2,a_3}\) and a set \(\bf B=a_3,a_4,a_5\). Then the union is:\[\bf A \cup B = A+B-(A ∩ B)= \left\{ a_1,a_2,a_3 \right\} + \left\{ a_3,a_4,a_5 \right\}-(\left\{ a_3 \right\})\]\[\bf \implies A \cup B = \left\{ a_1,a_2,a_3, a_3,a_4,a_5 \right\}-\left\{ a_3 \right\}=\left\{ a_1,a_2, a_3,a_4,a_5 \right\}\]You must evaluate your problem similarly. @AccidentalAiChan